OpenStudy (anonymous):

Imagine Bob, a shopkeeper, has a (square) tray of apples. While carrying the tray he trips and drops the apples all over the floor. His assistant comes out of the store room with a stack of 4 more trays. Clumsily he knocks the edge of the stack and a row of apples from each of the trays falls to the floor. Given that the total number of apples on the floor is 77 how many apples were on each tray to start?

OpenStudy (unklerhaukus):

the trays are 7x7 that is 49 apples per tray, adding 4x one row or 28 apples 49+4x7 =77

OpenStudy (anonymous):

well done, would you be kind to tell me how u got solving it?

OpenStudy (anonymous):

just explain to me how I can use the quadractic solving way to answer it, could you ,please

OpenStudy (anonymous):

And I'll give you another medal, two

OpenStudy (unklerhaukus):

well i kinda just took a rough guess and corrected it. knowing that the trays are square, i first guessed that the trays were 8x8 which is 64 , then adding 4x8 i got 96; to many next i tried a 7x7 square and found 49 + 7x4 =77 which is what the question demanded. there might be a simple algebra expression but i couldn't see it , so i just used guess and check method

OpenStudy (anonymous):

YOur cool man, thankx, brilliant, I was hoping more on a soliving a quadratic form by briliant , plain genious, thank you, you made my day

OpenStudy (unklerhaukus):

i guess you could have this equation \[x^2+4x=77\]\[x^2+4x-77=0\] and solve it using the quadratic formula

OpenStudy (anonymous):

Yup , brilliant , mate thanx

OpenStudy (unklerhaukus):

grin

OpenStudy (unklerhaukus):

yeah solving the quadratic gives us two solutions x=7 and x =-11, only one of these solutions makes sense in the real problem